# ADVANCING MATHEMATICS RESEARCH WITH GENERATIVE AI

### v2

@@ 10.1. Solving algorithmic research questions
In addition, generative AI models can generate code for Computer Algebra Systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory and other areas.
11. Examples of the use of generative AI models to address pattern-matching problems in higher mathematics
@@
12. Comparison of generative AI models
For the most difficult mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance. Though Gemini models performed slightly better in the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus and number theory in 2024.
Model MATH (%)
Claude 74.9
ChatGPT 76.6
Gemini 78.5
ChatGPT integrates features like WolframAlpha which can be enabled in the professional version. Another GPT called SciSpace has access to hundreds of millions of peer-reviewed journal papers.
13. General rules for writing prompts for advanced math questions:
@@
14. Structured prompts and system instructions
Structured prompts impose a logical framework on an LLM’s probabilistic generation process. The following examples of prompts have been shown to be useful in guiding generative AI models on how to approach complex mathematical questions [kojima2022]. These methods are also used internally in advanced LRM models like Gemini and ChatGPT.
14.1. Chain-of-thought reasoning
This refers to guiding the LLM to think step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. If the LLM shows its work, this reduces errors.
Example prompt:
Your task is to generate a detailed description of the representation of the Weyl group of $\mathfrak{sl}_{3}(\mathbb{C})$ on the dual space $\mathfrak{h}^{*}$ of its Cartan subalgebra $\mathfrak{h}$. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
[…]
14.2. Tree-of-thought prompt
This refers to guiding a generative AI model to explore multiple different deduction paths simultaneously such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
Example prompt:
is trivial or non-trivial. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
The emergence of LRMs and LCMs (Section 4) has altered the practice of prompt engineering.
15. Fact-checking: retrieval-augmented generation
@@ 16.1. Design your outputs
In generative AI models, there is a ‘system instruction’ feature which can be used to influence the type of output. You can enter a ‘pre-prompt’ with your own personal profile that governs the model’s behavior for the entire chat session.
@@ 17. Hybrid systems: generative AI models with Computer Algebra Systems or formal proof assistants
It is an active area of research to develop tools that link generative AI models with Computer Algebra Systems. One such tool is Google DeepMind’s AlphaGeometry. See also [KG2025].
Formal proof assistants like Lean allow for the expression and mechanical verification of mathematical proofs. Lean implements a version of Dependent Type Theory known as the Calculus of Inductive Constructions. This provides an alternative to the ZFC axioms of set theory. It allows the expression of mathematical claims, mechanically checks proofs of these claims and helps to find formal proofs.
Type Theory is advantageous for formalization because it embeds mathematical meaning directly into its syntax. Its system of Types prevents false statements and builds properties directly into the definition of its objects. In ZFC set theory, many of these logical statements require separate proofs.
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required. One of the benefits of the Lean environment is Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
@@
18. Should I collaborate with a generative AI model?
As we have discussed, generative AI, on its own, should not be viewed as an oracle or an authoritative source for mathematical output. Verification of output is a necessary part of using an AI model for mathematics. But it can be used as a creative collaborator that provides ideas, perspective and potential new directions. When combined with a Computer Algebra System, it can reliably perform laborious tasks and calculations. When paired with formal verification tools, it transforms the research process itself.
The author of this work now has several LLM+CAS+Lean-assisted research projects in infinite dimensional algebra and group theory. The use of generative AI models has been particularly revealing. There have been several instances where AI models have suggested novel research directions that had initially been dismissed as implausible or irrelevant. Yet, these suggestions turned out to yield fruitful and unexpected paths forward.

### v3

@@ 5.5. Using LLMs to generate Python code to run in SageMath
Generative AI models can generate code for Computer Algebra Systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory, ring theory and field theory.
Example prompt:
Generate Python code to find the minimum number of colors needed to color the Petersen Graph such that no two adjacent vertices share the same color (the ‘chromatic number’). Then create the Petersen Graph using the built-in SageMath generator. Find its chromatic number. Display the graph.
The emergence of LRMs and LCMs (Section 3.8) has altered the practice of prompt engineering.
5.6. Targeted prompts
@@
5.7. General rules for writing prompts for advanced math questions
@@ 6.1. Design your outputs
In generative AI models, there is a ‘system instruction’ feature which can be used to influence the output. Users can enter a ‘pre-prompt’ with their own personal profile that governs the model’s behavior for the entire chat session. Some models allow users to build AI agents that have persistent memory of settings. This capability allows for a type of ‘back-end’ engineering, where writing specific system prompts gives the user more direct control over the model’s performance. Utilizing this feature can significantly impact the amount of detail provided in the response and the level of rigor maintained in a mathematical proof.
@@ 6.2. Changing the ‘knobs’
6.3. Fact-checking: retrieval-augmented generation
7. Case Study: Combinatorial Group Theory
7.1. Why AI is a useful tool in Combinatorial Group Theory
As an example, we solve the word problem for the word $w=yxy^{-2}x^{-1}$:
We first rewrite $x^{-1}yx=y^{2}$ as $yx=xy^{2}$. Then
${yx}y^{-2}x^{-1}={xy^{2}}y^{-2}x^{-1}=xx^{-1}=1.$
Thus the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
7.2. Solving algorithmic research questions with AI
Algorithmic search problems in group theory, such as the word problem, conjugacy problem (deciding if two words are conjugate in a given group), and triviality problem (deciding if a group presentation presents the trivial group), are undecidable in general. Even in decidable cases, the search space for groups with many generators and intertwined relations is often astronomically large. The number of ways to apply relations to a word grows exponentially with the length of the word.
Gemini uses Chain-of-Thought processes for working with group presentations. The model does not just guess the final simplified word. It generates intermediate steps internally using a type of depth-first search where the model explores a simplification path. It can self-correct if the word complexity increases rather than decreases.
Gemini translates the group presentation $G=\langle X\mid R\rangle$ into a Python script using the sympy.combinatorics module.
  • It defines the free group on generators $X$.
  • It inputs the relations $R$.
  • It uses the library’s implementation of the Knuth-Bendix completion algorithm (or similar rewriting systems) to attempt to find a normal form for the word.
The following is an example of a prompt explicitly asking the model to follow Tree-of-Thought reasoning.
Example prompt:
is trivial or non-trivial. Proceed with the following tree of reasoning, evaluating each step before proceeding to the next:
@@ 8. Hybrid systems: generative AI models with Computer Algebra Systems or formal proof assistants
An active area of research is to develop tools that link generative AI models with Computer Algebra Systems [KG2025].
Formal proof assistants like Lean allow for the expression and mechanical verification of mathematical proofs. Lean implements a version of Dependent Type Theory known as the Calculus of Inductive Constructions (CIC). This is a constructive type theory, where in the Lean kernel (or core verifier) proofs are given explicitly, aligning with computation and the Curry–Howard correspondence. That is, every proof encodes an algorithm or a construction. However, in practice, Lean contains a model of ZFC set theory. It does not use ZFC axioms to check proofs, but with some additional axioms, it behaves like ZFC.
Type Theory is particularly advantageous for formalization because it embeds mathematical meaning directly into its syntax. Its system of Types prevents false statements and builds properties directly into the definition of its objects. In ZFC set theory, many of these logical statements require separate proofs.
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required.
Most users use Lean tactics. These are metaprograms that automate steps and give instructions on how to construct proofs. Meanwhile, Lean generates the rigorous low-level code for proof-checking.
One of the benefits of the Lean environment is Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
Lean Copilot is a framework that integrates LLMs into the Lean theorem prover to assist with formal proofs. It functions as an AI assistant by providing features like context-aware tactic suggestions. It selects relevant premises from Mathlib and searches for proofs.
@@
For example, Numina is an assistant that acts as a bridge between natural language mathematics proofs and Lean. It uses AI to generate formal proofs and uses Lean to verify their correctness.
9. Mathematician ++ generative AI collaboration
9.1. Human-LLM collaborative research
9.2. Why mathematicians should collaborate with generative AI models
As we have discussed, generative AI, on its own, should not be viewed as an oracle or an authoritative source for mathematical output. Verification of output is a necessary part of using an AI model for mathematics. But it can be used as a creative collaborator that provides ideas, perspective and potential new directions.
The most effective research strategy uses all available tools. When combined with a Computer Algebra System, AI can reliably perform laborious tasks and calculations. When paired with formal verification tools, it transforms the research process itself.
The author of this work now employs this collaborative framework in several LLM+CAS+Lean-assisted research projects in infinite dimensional algebra and group theory. The use of generative AI models has been particularly revealing in this context. There have been multiple instances where AI models have suggested novel research directions and ideas that had initially been overlooked or dismissed as implausible or irrelevant. Yet, these suggestions (once verified) turned out to yield unexpected paths forward.
Mathematicians could consider AI not merely as a search engine, but as a creative partner that can act as a catalyst for discovery.